英語 (English)

主題別英語 (Thematic English) (工 ((電 C)2 年))

鎌田スザーン · 非常勤講師 1 単位後期水 3 ・ 4

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 In today’s globalized society, the ability to communicate in English has become increasingly important to gain knowledge of world events and the ability to process technical information. In this class, students will participate in a variety of guided activities meant to foster communication skills. The emphasis will be on communication, as opposed to grammar.

【授業の概要】This class will be conducted entirely in English. You will be expected to speak, listen, read and write individually, in pairs and in groups. There will be a mid-term test and a final test at the end of the course.

【キーワード】 English communication 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】To become better at and more comfortable with using English for

communication. 【授業の計画】

1. Meeting and Greeting

2. Names

3. Kiss, Bow, or Shake Hands?

4. Prized Possessions

5. Cheers

6. What’s the Occasion?

7. Mid-term Test

8. First Dates

9. On the Job

10. A Gift for Me?

11. Feast on This

12. Looking Good

13. That’s Shocking!

14. Glued to the Tube

15. Review

16. Final Test 【教科書】 J-Talk: Conversation Across Cultures by Linda Lee, Kensaku Yoshida

and Steve Ziolkowski (Oxford University Press) 【参考書等】 [参考資料] 【成績評価の方法】 Grading will be based on attendance and participation (50%),

mid-term test ( 25%), final test (25%). 【再試験の有無】有【受講者へのメッセージ】 I hope that students will learn to enjoy using English.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221190 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 鎌田 . (オフィスアワー: suekamata@msn.com)

English

Thematic English (工 ((電 C)2 年))

名 Kamada · Part-time Lecturer 1 unit 後期水 3 ・ 4

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ In today’s globalized society, the ability to communicate in English has become increasingly important to gain knowledge of world events and the ability to process technical information. In this class, students will participate in a variety of guided activities meant to foster communication skills. The emphasis will be on communication, as opposed to grammar.

Outline⟩ This class will be conducted entirely in English. You will be expected to speak, listen, read and write individually, in pairs and in groups. There will be a mid-term test and a final test at the end of the course.

�������⟩ English communication Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ To become better at and more comfortable with using English for

communication.

Schedule⟩

1. Meeting and Greeting

2. Names

3. Kiss, Bow, or Shake Hands?

4. Prized Possessions

5. Cheers

6. What’s the Occasion?

7. Mid-term Test

8. First Dates

9. On the Job

10. A Gift for Me?

11. Feast on This

12. Looking Good

13. That’s Shocking!

14. Glued to the Tube

15. Review

16. Final Test Textbook⟩ J-Talk: Conversation Across Cultures by Linda Lee, Kensaku Yoshida

and Steve Ziolkowski (Oxford University Press) Reference⟩ [参考資料] Evaluation Criteria⟩ Grading will be based on attendance and participation (50%),

mid-term test ( 25%), final test (25%). Re-evaluation⟩ 有 Message⟩ I hope that students will learn to enjoy using English.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221190 Contact (Office-Hour, Room, E-mail)⟩

⇒ Kamada . (Office Hour: suekamata@msn.com)